Determine the molar solubility of copper(I) azide (CuN3CuN3) in a solution with a pH of 10.1210.12. Ignore activities. The ?sp for CuN3CuN3 is 4.9×10−94.9×10−9. The ?a for HN3HN3 is 2.2×10−5.
Answer #1

Let molar Solubility of CU M3 = xem CUM Cept + M3 x x ksp= {cut] [M] [cyt] [Mg] keq= THE UCH) 2.23*104 = 7.64101 - - x² = 16.pH = -log[ht] 16.12 = -log(H+3 => [ht] = 16-10-12-.586*1617 HNAN HI Horas a ko = [H] [Mg] a nos [NGH] Can Cut & Heag ksp = 4.